# A Word Order Synchronization Metric for Evaluating Simultaneous Interpretation and Translation

### v1

@@ Abstract
A Word Order Synchronization Metric for Evaluating Simultaneous Interpretation and Translation
Mana Makinae1, Katsuhito Sudoh1, Mararu Yamada2, Satoshi Nakamura1 |
@@ 4. Proposed Method
We propose a word order synchronization metric for SI and SiMT based on cross-lingual token alignment.
Suppose we have an English input “I ate apples yesterday.” and the corresponding Japanese translation {CJK}UTF8min “私は (I) / 昨日 (yesterday) / りんごを (apple) / 食べました。(ate),” where slashes represent boundaries of English words and Japanese chunks (in bunsetsu). Note that this is a simplified example and the actual tokenization is in subwords. If we have the correct alignment, we obtain the integer list [1, 4, 3, 2] to represent the word reordering happened in the translation. In this work, we use Spearman’s rank correlation coefficient instead of Kendall’s to take the reordering distance into account. In the example above, is equal to , which indicates little rank correlation.
Here, we suffer from many noisy alignments due to the non-parallelism mentioned above. To address this, we apply two heuristics to identify reliable bilingual alignment. First, we discard alignments with English function words because they do not always have clear word-level correspondence between different languages. We identify function words based on their part-of-speech tags using the function word list of SpaCy. Second, we ignore non-reliable alignments with cosine similarity below a certain threshold .
@@ 5. Experiments
We conducted the following two experiments to investigate the effectiveness of the proposed metric on actual SI and offline translation.
5.1. Setup
In the first experiment, we compared offline translation with SI using the proposed metric. We used NAIST-SIC-Aligned Zhao et al. (2023), a collection of English-to-Japanese SI for TED Talks with automatic segment-level alignment. It consists of triples of English talk transcripts, Japanese offline translation subtitles, and transcripts of Japanese SI by professional interpreters. We extracted 212 talks interpreted by interpreters with more than 15 years experience to focus on high-quality interpretations. SI by less experienced interpreters made it difficult to find semantic correspondence robustly due to interpretation errors. We reserve analyses of the less experienced interpretations as future work.
In the second experiment, we investigated the correlation of the proposed metric to human judgment. We used one talk session from Japan National Press Club (JNPC) Interpreting Corpus Matsushita et al. (2020b), a collection of English-to-Japanese SI for press conferences. The SI transcripts were evaluated by a Japanese-native professional interpreter using Multidimensional Quality Metrics (MQM) Lommel et al. (2014); Agarwal et al. (2023) from the viewpoint of her SI expertise. We chose 25 segments out of 172, which are longer than 30 words and have more than three aligned words, according to our findings in the first experiment.
5.2. Comparison between offline translation and simultaneous interpretation
@@
Finally, we investigated the relationship between word order synchronization and human judgment of the interpretation quality. Table 5 shows the Pearson correlation coefficients between the metrics scores (COMET-QE, BERTScore (F1), and the proposed metric) and the MQM error score. The proposed metric demonstrated a negative correlation to human judgment, while the others showed little correlation probably due to the difficulty of quality measurement of SI using the current automatic MT evaluation metrics. This result supports the usefulness of the measurement of word order synchronization for SI evaluation. The findings highlight the need for a specialized approach to the evaluation of SI and SiMT not in the word-by-word semantic equivalence but in the appropriate content delivery under real-time constraints.
@@ 6. Conclusion
In this work, we propose a word order synchronization metric based on rank correlation coefficients leveraging a simple word alignment motivated by BERTScore. Our metric revealed stronger word order synchronization of SI for long input segments than that of offline translation through the experiments using an English-to-Japanese SI corpus. Future work includes more robust synchronization evaluation for SI and SiMT with summarizations, paraphrasing, omissions, etc., and quality evaluation specialized in SI and SiMT.

### v3

# An Automatic Quality Metric for Evaluating Simultaneous Interpretation

@@ Abstract
The code is available at https://github.com/ahclab/Combined_Metrics.
@@
An Automatic Quality Metric for Evaluating Simultaneous Interpretation
Mana Makinae1, Katsuhito Sudoh1, Masaru Yamada2, Satoshi Nakamura1 |
@@ 4. Proposed Method
4.1. Synchro Metric
We propose a word order synchronization metric, abbreviated as the ‘synchro’ for SI and SiMT based on cross-lingual token alignment, enabling us to obtain the indices of maximum similarity among all tokens although the source and target languages are different.
Suppose we have an English input “I ate apples yesterday.” and the corresponding Japanese translation {CJK}UTF8min “私は (I) / 昨日 (yesterday) / りんごを (apple) / 食べました。(ate),” where slashes represent boundaries of English words and Japanese chunks (in bunsetsu). Note that this is a simplified example and the actual tokenization is in subwords. If we have the correct alignment, we obtain the integer list [1, 4, 3, 2] to represent the word reordering happened in the translation. In this work, we use Spearman’s rank correlation coefficient instead of Kendall’s to take the reordering distance into account. In the example, is equal to , which indicates little rank correlation.
Here, we suffer from many noisy alignments due to the non-parallelism mentioned above. To address this, we apply two heuristics to identify reliable bilingual alignment. First, we discard alignments with English function words because they have minimal lexical significance or convey grammatical connections among other words within a sentence, which is possible for a word or token present in the source to be absent in the target, or vice versa. We identify function words based on their part-of-speech tags using the function word list of SpaCy.
Second, we ignore non-reliable alignments with cosine similarity below a certain threshold . Multilingual BERT is suitable for getting token alignment on non-English language, however due to the limitation of Multilingual BERT performance and uniqueness in SI such as summarizations and omissions, low maximum similarity scores often end up in matching token incorrectly. To mitigate this concern, filtering with the words having high maximum similarity is necessary to accurately calculate rank correlation coefficient.
4.2. Combined Metric
We propose combined metrics of content words coverage and word order synchronization, abbreviated as the ‘combined’, as automatic simultaneous interpretation evaluation. Unlike synchro, this approach distinguishes itself through alignment method enhancements and the incorporation of content word coverage
To enhance its accuracy in calculation of rank correlation coefficients, we have improved word alignment method. In the first step, we tokenize a sentence into words using spaCy. Each word is then tokenized into subwords using the mBERT tokenizer. We proceed to obtain bilingual subword alignment using both Awesome Align and BERTScore. Subsequently, we extract subword alignment by identifying common alignments between Awesome Align and BERTScore, converting these subword alignments to word-level alignments. Based on the word alignment, we calculate word order synchronization by calculating Spearman’s rank correlation coefficient .
In addition, this metric introduces the coverage of content words. Our initial experiments of calculating word order synchronization among different SI ranks revealed that assessing word order synchronization is essential but not enough to capture the quality of the outputs, leading to the introduction of the content words coverage as an additional metric. We introduce content word coverage as follows to the idea from RIBES’s penalty. Content Words Coverage , where is the number of content words obtained in target outputs, and is the total number of content words in the source inputs.
@@ 5. Experiments
5.1. Comparison between SI and Offline Translation by Synchro Metric
5.1.1. Setup
We conducted the following two experiments to investigate the effectiveness of the synchro metric on actual SI and offline translation. In the first experiment, we compared offline translation with SI using the proposed metric. We used NAIST-SIC-Aligned Zhao et al. (2023), a collection of English-to-Japanese SI for TED Talks with automatic segment-level alignment. It consists of triples of English talk transcripts, Japanese offline translation subtitles, and transcripts of Japanese SI by professional interpreters. We extracted 212 talks interpreted by interpreters with more than 15 years experience to focus on high-quality interpretations. SI by less experienced interpreters made it difficult to find semantic correspondence robustly due to interpretation errors. We reserve analyses of the less experienced interpretations as future work.
In the second experiment, we investigated the correlation of the proposed metric to human judgment. We used one talk session from Japan National Press Club (JNPC) Interpreting Corpus Matsushita et al. (2020b), a collection of English-to-Japanese SI for press conferences. The SI transcripts were evaluated by a Japanese-native professional interpreter using Multidimensional Quality Metrics (MQM) Lommel et al. (2014); Agarwal et al. (2023) from the viewpoint of her SI expertise. This is based on our previous work when we hired a translation evaluator and SI evaluator to evaluate SI output with MQM, we observed different viewpoints between the two regarding accuracy and fluency. Particularly in fluency, the translation evaluator assigned negative scores to outputs if the target speech’s word order synchronized with the source speech, judging it as unnatural and damaging its fluency, whereas the SI evaluator did not share the same opinion. We consider these differences to be distinctive features specific to SI and asked evaluators to evaluate SI outputs with MQM, taking into account the unique characteristics of SI. We chose 25 segments out of 172, which are longer than 30 words and have more than three aligned words, according to our findings in the first experiment.
5.1.2. Analysis
@@
Finally, we investigated the relationship between word order synchronization and human judgment of the interpretation quality. Table 5 shows the Pearson correlation coefficients between the MQM error score and the metrics scores (COMET-QE, BERTScore (F1), and the proposed metric). The MQM is a less is better metric, meaning it is based on error analysis. A lower score indicates better quality, while a higher score reflects worse quality. In contrast, the proposed and other metrics follow a more is better metrics, where a higher score means better quality. When comparing error-based scores with accuracy-based scores, a negative correlation suggests that the two scoring methods are aligned, while a positive correlation indicates they are not. Using this interpretation, the proposed metric demonstrated a negative correlation to human judgment, while the others showed little correlation probably due to the difficulty of quality measurement of SI using the current automatic MT evaluation metrics. This result supports the usefulness of the measurement of word order synchronization for SI evaluation. The findings highlight the need for a specialized approach to the evaluation of SI and SiMT not in the word-by-word semantic equivalence but in the appropriate content delivery under real-time constraints.
Furthermore, we compare the efficacy of the MQM modified with SI manner to the original MQM, designed for evaluating translation outputs. Table 6 presents the Pearson correlation coefficients, revealing a stronger correlation with MQM modified with SI manner. Similar to the score interpretation in Table 5, Table 6 shows that the MQM modified with SI manner has a stronger negative correlation, suggesting that the modified version of MQM is more suitable for evaluating SI outputs than the original MQM, which was primarily designed for assessing translation quality.
5.2. Comparison among different SI ranks by Combined Metric
5.2.1. Setup
Metric | Score |
MQM | |
(original) | 0.132 |
MQM | |
(SI modified) | -0.497 |
Model | AER | Precision | Recall | F1 |
Best model (-entmax, multilingual ( = 0)) | 37.4 | 72.7 | 55.0 | 0.626 |
Model without fine-tuned softmax | 45.5 | 68.0 | 45.5 | 0.545 |
Combined | 52.3 | 73.5 | 35.3 | 0.477 |
Talk Name | A | B | S | Caption |
AlexanderWagner_2016X | 0.611 (0.433) | 0.799 (0.493) | 0.593 (0.480) | 0.465 (0.623) |
CharlesFleischer_2005 | 0.574 (0.438) | 0.637 (0.4644) | 0.679 (0.455) | 0.605 (0.617) |
ColinGrant_2012X | 0.543 (0.338) | 0.547 (0.470) | 0.614 (0.425) | 0.594 (0.587) |
DilipRatha_2014G | 0.482 (0.488) | 0.561 (0.526) | 0.575 (0.507) | 0.553 (0.604) |
HamishJolly_2013X | 0.473 (0.334) | 0.566 (0.328) | 0.526 (0.391) | 0.575 (0.664) |
IlonaSzabodeCarvalho_2014G | 0.328 (0.373) | 0.539 (0.417) | 0.583 (0.438) | 0.508 (0.646) |
JonathanDrori_2010U | 0.592 (0.461) | 0.541 (0.465) | 0.694 (0.485) | 0.657 (0.623) |
LaurelBraitman_2014S | 0.425 (0.360) | 0.566 (0.3932) | 0.548 (0.401) | 0.586 (0.600) |
MiruKim_2008P | 0.517 (0.437) | 0.521 (0.428) | 0.412 (0.470) | 0.412 (0.668) |
NickBostrom_2015 | 0.487 (0.358) | 0.652 (0.375) | 0.675 (0.385) | 0.517 (0.611) |
ShlomoBenartzi_2011S | 0.422 (0.433) | 0.477 (0.422) | 0.502 (0.448) | 0.605 (0.630) |
ShubhenduSharma_2016S | 0.498 (0.444) | 0.663 (0.507) | 0.415 (0.487) | 0.610 (0.626) |
TaraWinkler_2016X | 0.420 (0.444) | 0.780 (0.483) | 0.595 (0.408) | 0.568 (0.613) |
Source | And in the case of drugs in order to undermine this fear and prejudice that surrounds the issue we managed to gather and present data that shows that today’s drug policies cause much more harm than drug use per se and people are starting to get it . |
Target | {CJK}UTF8minドラッグに関して恐怖であったり偏見をなくして行く為にデータを見せて行く今の政策の方がドラッグよりもどんどん 悪い結果になっているそしてそれが人々には分かり始めました。 |
Aligned Words | cause, harm, drugs, drug, people, fear, starting, prejudice, data, today, policies |
Synchro | 0.9818 |
Coverage | 0.478 |
Talk Name | SI rank | MQM (SI-modified) | Combined |
IlonaSzabodeCarvalho_2014G | B | 238 | 0.225 |
IlonaSzabodeCarvalho_2014G | S | 92 | 0.255 |
TaraWinkler_2016X | B | 10 | 0.342 |
TaraWinkler_2016X | S | 105 | 0.243 |
We conducted the following two experiments to investigate the combined metric’s effectiveness among different SI ranks. In the first experiment, we used the Kyoto Free Translation Task (KFTT) Neubig (2011), a task for Japanese-English translation that focuses on Wikipedia articles related to Kyoto to investigate the effectiveness of the improved word alignment method. Among the datasets, we extracted test data from Alignment Data.
In the second experiment, we used NAIST-SIC-Aligned Zhao et al. (2023), a collection of English-to-Japanese SI for TED Talks with automatic segment-level alignment to investigate a combined metric’s effectiveness. It consists of triples of English talk transcripts, Japanese offline translation subtitles, and transcripts of Japanese SI by professional interpreters. We extracted 13 talks because these talks are only talks translated by all different experienced interpreters(S, A, and B).
The process of human judgment follows the same as in the first metrics’ proposal.
5.2.2. Analysis
In the first experiment, Table 7 illustrates the effectiveness of the improved word alignment method. Despite the lowest values in AER, Recall, and F1, our method shows the highest Precision. This suggests that when the proposed model predicts a positive instance, it tends to be accurate. However, it also indicates a notable omission of actual positive instances. The further analysis of its word alignment is in Appendices 9.1.
Our proposed method, though providing fewer alignments, maintains a relatively high accuracy, making it suitable for this task. Therefore, we use this word alignment method for the rest of the research.
In the second experiment, Table 8 presents the word order synchronization and content words coverage for each talk across various SI ranks and Captions. For word order synchronization, there is a relative synchronization between source inputs and target outputs in SIs, better than in Caption. As for the content words coverage, Caption shows the highest coverage, which is an anticipated finding. However, what is noteworthy is that in certain talks, the coverage is higher in the B rank than in the S rank. This observation is surprising, as we initially hypothesized that higher ranks would correspond to better quality.
To further investigate this, we provide a detailed analysis using an example shown in Table 9. Our examination resulting from this experiment reveals that regardless of the rank of the SI, if a sentence achieves both high synchronicity and coverage, the outputs tend to demonstrate high quality.
Finally, we examine whether the results obtained from our second proposal correlate with human judgment. In this analysis, we chose two talks IIlonaSzabodeCarvalho_2014G and TaraWinkler_2016X with a focus on S rank and B rank interpreters’ outputs. This choice was based on the observation that, contrary to our initial intuition, in one talk, S rank interpreters performed better, while in the other, B rank interpreters exhibited superior output.
In Table 8, the combined metric in IIlonaSzabodeCarvalho_2014G reveals that the S rank outperforms the B rank. Specifically, the synchronization score and content word coverage for the S rank are 0.583 and 0.438, respectively, while those for the B rank are 0.539 and 0.417. In terms of human evaluation scores, the same tendency is observed in Table 10, with penalty scores for S rank interpreters at 92, as opposed to 238 for B rank interpreters. Consequently, both quality evaluation reveals that the output of S rank interpreters is superior to that of B rank interpreters.
On the other hand in TaraWinkler_2016X combined metric indicates that B rank surpasses S rank. Notably, the B rank achieves a synchronization score of 0.780 and a content words coverage of 0.438, while the S rank scores 0.595 and 0.408 in the same metrics. Human evaluation scores represent the same trend, with penalty scores of 10 for B rank and 105 for S rank interpreters, showing the superiority of B rank outputs. This result indicates that, despite having the most experience, the S rank doesn’t consistently deliver the best performance.
@@ 6. Conclusion
In this work, we proposed a word order synchronization metric, abbreviated as the ’Synchro metric’, employing BERTScore to obtain token alignments directly from the source and target. We also introduced combined metrics of content words coverage and word order synchronization, abbreviared as combined metric, leveraged alignment information from Awesome Align and combined it with BERTScore to obtain reliable token alignments.
Our first metric revealed stronger word order synchronization of SI for long input segments than that of offline translation through the experiments using an English-to-Japanese SI corpus.
Our second metric demonstrated that outputs achieving both high word order synchronization and high coverage of content words are a sign of high-quality outputs in some cases. It also revealed high word order synchronization and high content words coverage are not enough for the high quality outputs in some cases when the significant information is omitted and hurts overall understanding.
Word order synchronization between the source and target has not received sufficient attention in both SI and SiMT, despite its crucial role solving the trade-off between quality and latency. Our focus on these issues led to the proposed token and word semantic-based evaluation, recognizing that word order gaps significantly impact latency in both SI and SiMT. Future work includes more robust synchronization evaluation for SI and SiMT with summarizations, paraphrasing, omissions, etc., and quality evaluation specialized in SI and SiMT.
@@ 7. Bibliographical References
- Neubig (2011) Graham Neubig. 2011. The kyoto free translation task (kftt).
@@ 8. Language Resource References
9. Appendices
9.1. Additional Word Alignment
Sentence Pair | {CJK}UTF8min東西 線 の 駅 は 駅 ごと に ステーション カラー が 制定 さ れ て い る が 、 三条 京阪 駅 の ステーション カラー は 牡丹 色 。 |
each station on the tozai line has its own color , and the color used at sanjyo-keihan station is scarlet tinged with purple . | |
Human-annotated | {CJK}UTF8mineach:ごと, each:に, station:駅, tozai:東西 |
Word Alignment | {CJK}UTF8minline:線, has:制定, has:さ, has:れ, has:て, has:い, has:る, its:ステーション, its:カラー, own:ステーション, own:カラー, color:ステーション, color:カラー, ,:、, color:ステーション, color:カラー, at:の, sanjyo-keihan:三条, sanjyo-keihan:京阪, station:駅, is:は, scarlet:牡丹, scarlet:色, tinged:牡丹, tinged:色, with:牡丹, with:色, purple:牡丹, purple:色, .:。 |
Word Alignment | {CJK}UTF8minstation:駅, line:線, has:は, color:カラー |
from Combined | {CJK}UTF8minand:、, color:カラー, at:の, sanjyo-keihan:三条, sanjyo-keihan:京阪, station:駅, is:は, scarlet:牡丹, purple:色 |
Source | But such ardent watchfulness can lead to anxiety so much so that years later when I was investigating why so many young black men were diagnosed with schizophrenia six times more than they ought to be I was not surprised to hear the psychiatrist say Black people are schooled in paranoia. |
Target | {CJK}UTF8min不安感がありましたそれで何年か経って何で若い黒人が統合失調症になる可能性が六倍あると言う事が分かりました個人黒人と言うのはパラノイアだと言う風に言われてるんです。 |
Combined | {CJK}UTF8minwatchfulness:感, anxiety:不安, years:年, young:若い, black:黒人, men:黒人, diagnosed:症, schizophrenia 失調, schizophrenia:症, times:倍, paranoia:パラノイア |
Awesome | {CJK}UTF8minwatchfulness:感, lead:あり, anxiety:不安, years:年, later:で |
Align | {CJK}UTF8minyoung:若い, black:黒人, men:黒人, diagnosed:症, schizophrenia:失調, schizophrenia:症, times:倍, hear:言う, Black:黒人, schooled:言わ, paranoia:パラノイア |
BERTScore | {CJK}UTF8minardent:感, watchfulness:感, lead:感, anxiety:不安, anxiety:感, years:年, investigating:統合, investigating:が, young:若い, black:黒人, men:黒人, diagnosed:症, schizophrenia:失調, schizophrenia:症, times:倍, ought:が, surprised:言わ, surprised:に, hear:言わ, psychiatrist:パラノイア, psychiatrist:黒人, psychiatrist:症, Black:黒人 , people:黒人, schooled:統合, schooled:パラノイア, paranoia:パラノイア |
To investigate the proposed word alignment further, we examine word alignment through a specific example in Table 11. The word alignment results obtained correctly by the improved word alignment method involve content words such as station, line, and color. Conversely, instances where the proposed method fails to obtain word alignments are characterized by function words such as each, own and is. Given that the aim of this research is to investigate word order synchronization, with a focus on content words rather than function words and not specifically on enhancing word alignment, the observed outcome of the highest precision is enough for this study.
Additionally, Table 12 shows an example of word alignment method comparison in a real scenario where source inputs and target outputs aren’t in an equal relationship due to specific characteristics in SI such as omission, summarization etc.. We compared its effectiveness with a previous method and awesome align. The previous method, using BERTScore’s greedy match, often aligns words incorrectly, leading to low accuracy. Awesome align also faces misalignment issues, which can negatively impact word order distance calculations, which is the focus of our research.
9.2. Detailed Analyses with Combined Metric
Segment Length | Combined | BERTScore (F1) |
All | -0.130 | -0.049 |
0.094 | 0.011 | |
-0.183 | 0.035 | |
-0.225 | 0.113 | |
-0.101 | 0.077 | |
-0.262 | 0.105 |
For a detailed analysis, we examined the rank correlation to human judgment between the proposed method and BERTScore (F1) in each sentence and the Table 13 shows its details.
In the row ’All’ in Table 13, a very weak correlation is observed more in combined metric than in BERTScore. From the findings in previous results and analysis, which highlights the significant advantage of the combined metric as the segment length increased, we further breakdown its results based on varying segment lengths. In the proposed method, we observed a slight correlation improvement as the segment length increased, whereas in BERTScore (F1), no correlation was observed regardless of the segment length. These results indicate that our combined metric excels at evaluating relatively long sentences, though some improvement is needed when assessing short segments.
Source | Each and every one of us has the power to change the world. |
Target | {CJK}UTF8min私達一人一人は、世界を変える力を持っているのです。 |
Combined | -0.666 (-1 (Synchro) * 0.666 (Coverage)) |
MQM | 0 |
Table 14 shows the challenge combined metric faces when evaluating short segments, which rooted in the interaction between the two components: the multiplication of content word coverage and the word order synchronization score. Despite a higher coverage of words, a combined metric evaluates short segments unfairly due to the unavoidable challenge of low word order synchronization in short segments, where word reordering occurs naturally even in SI. In long segments, a combined metric aligns with human evaluation results in some cases, while in others, it does not.
Source | What I ’m talking about and what the Global Commission advocates for is creating a highly regulated market where different drugs would have different degrees of regulation. |
Target | {CJK}UTF8min私が言いたいと思ってるのは非常に規制の高いマーケットを作ると言う事ですドラッグのレベルによって規制のレベルを変えると言う事です。 |
Combined | 0.417 (0.893 (Synchro)) * 0.467 (Coverage) |
MQM | 0 |
Source | Civil society diplomats do three things They voice the concerns of the people are not pinned down by national interests and influence change through citizen networks not only state ones. |
Target | {CJK}UTF8min人々の声を取り上げ国の声では無くて人々の声を取り上げて来るのです。 |
Combined | -0.09 (-0.5 (Synchro) * 0.188 (Coverage)) |
MQM | 10 |
The following is the example in the long segment when a combined metric aligns with human judgment. In Table 15, the sentence is not penalized in human evaluation, indicating the quality of its output is high. In a combined metric, the multiplication of the average synchronization score and content words coverage, resulting in 0.255, indicating that the multiplied score of 0.417 expresses higher quality compared to the average score. In Table 16, the evaluator penalizes this sentence as 10, and combined metric gives -0.009). In such instances where obvious omission happens, the combined metric could capture its tendency due to the calculation of its content words.
Source | In just a few years, we not only changed national legislation that made it much more difficult for civilians to buy a gun, but we collected and destroyed almost half a million weapons. |
Target | {CJK}UTF8minこの何年の内にもっともっと市民が銃を買うのを難しくないといけない。そして半分の銃は捨て去って行かないといけない。 |
Combined | 0.612 (0.883 (Synchro) * 0.692 (Coverage)) |
MQM | 10 |
Table 17 shows a examples when combined metric cannot fairly evaluate long segments. In the example, the evaluator penalizes this sentence as 10, while combined metric gives 0.612. In such instances, the translations may initially seem coherent due to an adequate number of target outputs corresponding to the source inputs. However, more detailed examination reveals the lack of connection between individual words results in outputs that are challenging to understand. A combined metric cannot follow the connectivity of each word since we only count the content words coverage and the word order synchronization, leading to discrepancy between the human judgment and proposed method.
Source | Civil society diplomats do three things They voice the concerns of the people are not pinned down by national interests and influence change through citizen networks not only state ones. |
Target | {CJK}UTF8minこの外交官の仕事は、三つがあります。一つは、まず人の声を代弁する。そして、人に対して影響を与えていくと言う事です。 |
Combined | 0.353 (0.616 (Synchro)) * 0.375 (Coverage)) |
MQM | 10 |
In Table 18, evaluator penalizes this sentence as 10, while combined metric gives 0.353. Critical omission is also difficult to detect in such cases when an adequate number of target outputs corresponding to the source inputs is observed at first glance. This drawback could potentially be solved by assigning weights based on the importance of content words, instead of treating all words equally in the current proposed method. While we have demonstrated a significant advantage of combined metric over BERTScore, there is room for further improvement in both short and long segments. We reserve further enhancement of our automatic evaluation for more precise results as future work.
9.3. Comparison between SI and Offline Translation using word alignment method in 4.2
Additionally, we investigated the validity of method 5.2 under the assumption that alignment method in 5.2 outperforms 4.1. Given the previous findings that our method excels at evaluating long segments, we chose the same dataset in the second experiment in 5.1.1 because each segment in JNPC is longer than TED Talks.
Segment Length | SI | Translation |
All | 0.686 | 0.455 |
0.479 | 0.011 | |
0.736 | 0.512 | |
0.786 | 0.563 | |
0.790 | 0.596 | |
0.885 | 0.740 |
We conducted two experiments following the process in 5.1.1. In the first experiment, we compared the word order synchronization using combined metric’s word alignment. The results are presented in Table 19. As expected, across the entire dataset, SI synchronizes more closely with the source speech than translation, and this tendency becomes more evident as segments get longer.
Metric | Score |
COMET-QE | 0.161 |
BERTScore (F1) | 0.1111 |
Synchro | -0.497 |
Combined (synchro only) | -0.011 |
Combined | 0.022 |
In the second experiment, we compared synchro score alone and combined metric with human evaluation Table 20. However, neither the synchro metric nor the combined metric showed any correlation with human judgment.
Source | And if we are to have a union that would be stable and that would be at peace because all the peoples accept that it is a union that will guarantee their security then we will have to amend the constitution and we have to work towards it. |
Target | {CJK}UTF8min我々は作りたい国が安定してそしてすべての国民が彼らの違う安全を保障してくれると信じてくれればそれならば憲法を改正できると思っております。 |
Combined | 0.573 (0.9 (Synchro) * 0.636363636 (Coverage)) |
MQM | 79.8 |
Comments by | {CJK}UTF8min「安定して平和な連邦国家を築くのであれば」が大部分抜け, |
SI Evaluator | {CJK}UTF8min(誤訳)→「すべての国民が自分達の安全を保障してくれるのは連邦国家であると受け入れるのであれば」, |
{CJK}UTF8min(誤訳)→「憲法を改正しなくてはならず、そのために努力しなければならない」の意 |
This results might follows the same tendency observed in Table 17 and Table 18 indicating that critical mistranslation and omission are difficult to detect when an adequate number of target outputs corresponding to the source inputs is observed at first glance (Table 21). This highlights a current limitation in a combined metric, which treats each content word equally. However, the positive outcome is that, owing to the improved alignment ability in combined metric, we successfully extracted 48 sentences meeting the criteria of segment lengths longer than 30 words and containing more than three aligned words. This represents an improvement compared to the synchro metric, which could only extract 25 sentences under the same conditions.